# Forecasting CPI inflation under economic policy and geo-political uncertainties

### v1

@@ Abstract
Forecasting a key macroeconomic variable, consumer price index (CPI) inflation, for BRIC countries using economic policy uncertainty and geopolitical risk is a difficult proposition for policymakers at the central banks. This study proposes a novel filtered ensemble wavelet neural network (FEWNet) that can produce reliable long-term forecasts for CPI inflation. The proposal applies a maximum overlapping discrete wavelet transform to the CPI inflation series to obtain high-frequency and low-frequency signals. All the wavelet-transformed series and filtered exogenous variables are fed into downstream autoregressive neural networks to make the final ensemble forecast. Theoretically, we show that FEWNet reduces the empirical risk compared to single, fully connected neural networks. We also demonstrate that the rolling-window real-time forecasts obtained from the proposed algorithm are significantly more accurate than benchmark forecasting methods. Additionally, we use conformal prediction intervals to quantify the uncertainty associated with the forecasts generated by the proposed approach. The excellent performance of FEWNet can be attributed to its capacity to effectively capture non-linearities and long-range dependencies in the data through its adaptable architecture.
@@ 1 Introduction
The monetary policy framework at central banks relies heavily on forecasts of various macroeconomic variables. Forecasts have also grown in importance as a means of communication for central banks during the past 20 years, and they have the potential to shape public and market expectations. Numerous studies have examined central bank forecasting from various angles due to the paramount significance of these predictions[1, 2, 3, 4, 5]. Forecasting, particularly in the context of different macroeconomic policy variables, is one of the major focus areas in this domain. Inflation is one such policy variable considered instrumental in designing an economy’s overall monetary policy. Thus, a deeper understanding of the factors that help forecast inflation accurately is fundamental to policymakers at the central bank [1]. Over the past few decades, emerging market economies such as Brazil, Russia, India, and China (BRIC) have undergone significant changes in their macroeconomic conditions, including shifts in monetary policy frameworks such as inflation targeting, increased globalization, and the impact of events like the financial crisis and the recent pandemic. These shifts make predicting key macroeconomic variables like consumer price index (CPI) inflation quite challenging [2]. Given these dynamic changes in the macroeconomic environment, factors like economic policy uncertainty and geopolitical risk might offer critical insights to forecast inflation accurately.
@@
Due to the significance of inflation forecasting exercise for monetary policy, central banks often depend on distinct models, namely structural and non-structural. The structural models, which are different versions of Phillips curve models, are derived from economic theory. On the other hand, the non-structural models, such as linear univariate and multivariate time-series models, aim to utilize the unique characteristics of the data without explicitly depending on any economic theory [3]. Non-structural models often prioritize forecast accuracy above causal inference and are particularly valuable for short-term, out-of-sample forecasting. Linear time series models such as autoregressive integrated moving average (ARIMA) and vector autoregressive (VAR) are limited in their ability to account for business cycles, periods of extreme volatility caused by macroeconomic uncertainties, and structural changes. As a result, non-linear, hybrid, and machine learning models have become increasingly popular to address these shortcomings. Furthermore, the existing linear time series forecasting methods bring several undesirable properties, ranging from high sensitivity to model specification to high-frequency data requirements [4]. This is especially important for forecasting inflation or similar macroeconomic variables that are only available at the monthly, quarterly, and annual frequency. This issue is more serious for emerging market economies where the quality of macroeconomic data is not up to the mark. Given such issues related to inflation data in emerging countries, machine learning (ML) models and combined approaches provide an opportunity to improve forecast performance.
@@
Recently, an increasing number of studies have been conducted on applying ML models for inflation forecasting, both for advanced and emerging market economies. A study conducted in the United States utilized a basic feed-forward neural network to forecast quarterly CPI inflation. The study concluded that this approach was successful in predicting CPI inflation [5]. Neural networks proved more effective than traditional models in predicting the monthly inflation rate for the OECD countries [6]. Their research revealed that neural network models outperformed simple AR models in 45% of the nations, on average, whereas simple AR models performed better in 23% of the countries. Recurrent neural networks, in their various forms, have been discovered to be beneficial for predicting inflation in developed countries [7, 8, 9, 10]. In addition, a recent study compared various machine learning models, including lasso regression, random forests, and deep neural networks, to analyze inflation in the United States. The study also considered several external factors, such as cash and credit availability, online prices, housing prices, consumer data, exchange rates, and interest rates [11]. Their research has shown that machine learning models with many covariates consistently outperformed conventional techniques. This superiority can be attributed to nonlinear relationships between historical macroeconomic variables and inflation. This has dispelled doubts about using machine learning and contemporary deep learning architectures in analyzing big data in economics [12]. For emerging market economies, the study of inflation forecasting using advanced ML models is in a nascent stage. Applications of neural networks are evident in developing countries like India, Brazil [13], Russia [14], Turkey, South Africa [15], and China [16]. However, none of the above studies, to the best of our knowledge, have considered exogenous factors like macroeconomic uncertainties for inflation forecasting, especially for emerging markets.
@@
Following events such as the global financial crisis, it has been observed that many countries worldwide have encountered heightened uncertainty. This has had an impact on the decision-making process of individuals and consequently on the overall macroeconomy [17, 18, 19]. Although previous studies have primarily examined the connection between uncertainty and the overall economy, the evaluation of the changing relationship between measures of uncertainty and inflation is still in its early stages. It has not yet yielded a definitive conclusion[20, 21]. Gaining insight into the intricacies of inflation dynamics and uncertainty metrics is quite critical for policymakers. Multiple recent studies indicate that both economic policy uncertainty and geopolitical risk significantly influence the movement of inflation within a given sample [22, 20, 23, 24, 25, 26]. Consequently, inflation forecasting models that do not incorporate these uncertainties are deemed to be incorrectly specified. The study conducted by [23] examines the effectiveness of different uncertainty measures in predicting inflation. In particular, they use US inflation data and the economic policy uncertainty index (EPU) to compare how well the vector autoregressive fractionally integrated moving average (VARFIMA) model performs at making predictions compared to other models. The study conducted by [27] examined the predictive impact of Economic Policy Uncertainty (EPU) on stock market volatility in the United States and the United Kingdom. The findings suggest that EPU has a significant positive effect on volatility. However, this positive relationship weakens when the signal quality is low. [28] asserted that Economic Policy Uncertainty (EPU) has a crucial impact on enhancing the predictive accuracy of models forecasting stock market volatility. The study conducted by [29] investigated the impact of policy uncertainty in the United States and Europe on the short-term pricing model for gold. The study conducted by [30] investigated the impact of Economic Policy Uncertainty (EPU) on predicting future real economic activities in the UK and European economies. For an extensive analysis and thorough examination of different methods for quantifying uncertainty, as well as the impact of economic policy uncertainty on various economic activities such as stock market returns, corporate capital investments and spending, corporate finance, and risk management, please refer to the publication by [31]. On the other hand, the effect of geopolitical risk uncertainty (GPRC) has been studied extensively by [24]. A recent study by [32] aims to link geopolitical risk with the reduction of firm values as well. The studies covering the impact of various uncertainty measures on different economic activities are still in the nascent phase, and this study in turn, presents a way to accommodate the roles of EPU and GPRC in forecasting inflation numbers.
@@
Wavelet analysis has shown remarkable promise in analyzing financial time series, due to its ability to distillate crucial information from the financial data. It is their structural flexibility that allows them to handle irregular time series data more effectively for long-term time series forecasting [33]. Its ability to spot the time regime shifts and breakpoints allows it to discover important signals within the time series data. Broadly, wavelets are functions that can decompose the original time series into two sub-components: one concerning the low-frequency or deterministic component and the other concerning the high-frequency or stochastic component of the series. This process aims to reduce the effect of ‘noise’ in the forecast and is found to be quite useful in dealing with a variety of non-stationary signals, a common feature of financial and macroeconomic time series data, as they are constructed over finite time intervals and can represent the ‘local’ elements in both time and scale [34]. Several applications of wavelet analysis can be found in forecasting exchange rates, GDP growth, crude oil prices, stock market prices, Taiwan stock index, S&P dividend yield, trade prices, expenditure and income, price movements, money growth, and inflation [35, 36], volatility in foreign exchange markets, car sales from the last decades [37]. A wide variety of hybrid models combining wavelets and neural networks have been studied for forecasting natural gas prices [38], future trading [39], short-term electricity load [40, 41], energy price forecasting [42], oil prices [35], stock index [43]. Apart from forecasting, the application of wavelet analysis can be found in various areas of finance: [44] analyzed the association between stock price and inflation rate using wavelet analysis, [45] explored the aspect of assessment of the risk in the emerging markets, [46] studied the value at risk in the metal markets. [47] examined the impact of the global financial crisis on the multi-horizon nature of systematic risk and market risk using a wavelet multi-scale approach. [48] used wavelet neural networks for the prediction of corporate financial distress. However, none of the above studies have utilized inflation series and its causal counterparts for long-term forecasting.
@@
Against this background, the primary aim of this study is to create a filtered ensemble wavelet neural network (FEWNet) model. This model belongs to a novel category of machine learning models that combine traditional neural networks with wavelet models. The objective is to generate accurate long-term forecasts for CPI inflation in BRIC countries while taking into account economic and geo-political uncertainties. FEWNet can produce reliable long-term forecasts for CPI inflation of BRIC countries conditioned on economic and geopolitical uncertainties. In analyzing macroeconomic policy variables like CPI inflation, the wavelet decomposition can extract the low-frequency components of the series without sacrificing the fundamental characteristics of the original time series. This model also integrates a layer for feature engineering aspects by extracting the trend components using the Hodrick-Prescott (HP) filter and cyclical elements using the Christiano-Fitzgerald (CF) ideal band-pass filter in the forecasting exercise. Initially, the observed time series (CPI inflation) and exogenous factors (EPU and GPRC) undergo decomposition into their trend and cyclical components using HP and CF filters in the first phase. Subsequently, in the second phase, all the trend and cyclical components are modeled as auxiliary inputs within the FEWNet framework, along with the wavelet-transformed details and smooth coefficients of the CPI inflation. This comprehensive approach aims to generate forecasts for 12 months and 24 months ahead. Conditioning inflation forecasts on policy uncertainty and risk uncertainty would capture the inflation or deflation effects of uncertainty. We argue using wavelet coherence analysis that there exists a relationship between EPU and GPRC with that of CPI inflation for BRIC economies. We show that the data-generating properties of inflation (persistence and volatility) are associated with the movements in EPU and GPRC. Proposed FEWNet generates out-of-sample forecasts in a recursive manner, and using several statistical metrics and tests, we compare its performance with alternative forecasting methods. Theoretical results on empirical risk minimization are also presented for the FEWNet. Apart from improvement in forecast accuracy measures of CPI inflation for the BRIC countries, another critical area that interests the policymakers is the knowledge of how various forms of structural uncertainties can influence the movement or complex dynamic behavior of different macroeconomic variables. The performance of FEWNet is superior to its counterparts in all three economies, except for China. This demonstrates that FEWNet is highly effective in handling complicated, non-stationary time series in forecasting tasks.
@@
The paper is organized in the following way. Section 2 explains the macroeconomic series considered in this study, various data pre-processing strategies, and the causal association between the uncertainty indices and CPI inflation. Section 3 focuses on the proposed methodology with a detailed overview of its architectural design and Section 4 theoretically establishes the robustness of the method from an empirical risk minimization perspective. Section 5 discusses the experimental evaluation and the process of deriving a conformal prediction interval for the FEWNet framework. Finally, Section 6 ultimately wraps up the study by presenting a thorough analysis of the main discoveries and outlining potential future directions for the research.

### v2

# Forecasting CPI inflation under economic policy and geopolitical uncertainties

@@ Abstract
Forecasting consumer price index (CPI) inflation is of paramount importance for both academics and policymakers at the central banks. This study introduces a filtered ensemble wavelet neural network (FEWNet) to forecast CPI inflation, which is tested on BRIC countries. FEWNet breaks down inflation data into high and low-frequency components using wavelets and utilizes them along with other economic factors (economic policy uncertainty and geopolitical risk) to produce forecasts. All the wavelet-transformed series and filtered exogenous variables are fed into downstream autoregressive neural networks to make the final ensemble forecast. Theoretically, we show that FEWNet reduces the empirical risk compared to fully connected autoregressive neural networks. FEWNet is more accurate than other forecasting methods and can also estimate the uncertainty in its predictions due to its capacity to effectively capture non-linearities and long-range dependencies in the data through its adaptable architecture. This makes FEWNet a valuable tool for central banks to manage inflation.
keywords:
Inflation Forecasting , Wavelets , Neural Networks , Empirical Risk Minimization , Conformal Prediction Intervals
[…]
@@ 1 Introduction
The monetary policy framework at central banks relies heavily on forecasts of various macroeconomic variables. Forecasts have also grown in importance as a means of communication for central banks during the past 20 years, and they have the potential to shape public and market expectations. Numerous studies have examined central bank forecasting from various angles due to the paramount significance of these predictions Stock & Watson (2007); Pratap & Sengupta (2019); Faust & Wright (2013); Smalter Hall & Cook (2017); Nakamura (2005). Forecasting, particularly in the context of different macroeconomic policy variables, is one of the major focus areas in this domain. Inflation is one such policy variable considered instrumental in designing an economy’s overall monetary policy. Thus, a deeper understanding of the factors that help forecast inflation accurately is fundamental to policymakers at the central bank Stock & Watson (2007). Over the past few decades, emerging market economies such as Brazil, Russia, India, and China (BRIC) have undergone significant changes in their macroeconomic conditions, including shifts in monetary policy frameworks such as inflation targeting, increased globalization, and the impact of events like the financial crisis and the recent pandemic. These shifts make predicting key macroeconomic variables like consumer price index (CPI) inflation quite challenging Pratap & Sengupta (2019). Given these dynamic changes in the macroeconomic environment, factors like economic policy uncertainty and geopolitical risk might offer critical insights to forecast inflation accurately.
@@
Due to the significance of inflation forecasting exercise for monetary policy, central banks often depend on distinct models, namely structural and non-structural models. The structural models, which are different versions of Phillips curve models, are derived from economic theory. On the other hand, the non-structural models, such as univariate and multivariate time-series models, aim to utilize the unique characteristics of the data without explicitly depending on any economic theory Faust & Wright (2013). Non-structural models often prioritize forecast accuracy above causal inference and are particularly valuable for short-term out-of-sample forecasting. Linear time series models such as autoregressive integrated moving average (ARIMA) and vector autoregressive (VAR) are limited in their ability to account for business cycles, periods of extreme volatility caused by macroeconomic uncertainties, and structural changes. As a result, non-linear, hybrid, and machine learning models have become increasingly popular in addressing these shortcomings. Furthermore, the existing linear time series forecasting methods bring several undesirable properties, ranging from high sensitivity to model specification to high-frequency data requirements Smalter Hall & Cook (2017). This is especially important for forecasting inflation or similar macroeconomic variables that are only available at the monthly, quarterly, and annual frequency. Machine learning (ML) models and combined approaches provide an opportunity to improve the forecast performance of the inflation data in emerging economies.
@@
Recently, an increasing number of studies have been conducted on applying ML models for inflation forecasting, both for advanced and emerging market economies. A study conducted in the United States utilized a basic feed-forward neural network to forecast quarterly CPI inflation. The study concluded that this approach was successful in predicting CPI inflation Nakamura (2005). Neural networks proved more effective than traditional models in predicting the monthly inflation rate for the OECD countries Choudhary & Haider (2012). Their research revealed that neural network models outperformed simple AR models in 45% of the nations, on average, whereas simple AR models performed better in 23% of the countries. Recurrent neural networks, in their various forms, have been useful for predicting inflation in developed countries Chen et al. (2001); McAdam & McNelis (2005); Almosova & Andresen (2023); Barkan et al. (2023). In addition, a recent study compared various machine learning models, including lasso regression, random forests, and deep neural networks, to analyze inflation in the United States. Medeiros et al. (2021) also considered several external factors, such as cash and credit availability, online prices, housing prices, consumer data, exchange rates, and interest rates Medeiros et al. (2021). Their research has shown that machine learning models with many covariates consistently outperformed conventional techniques. This superiority can be attributed to non-linear relationships between historical macroeconomic variables and inflation. This has dispelled doubts about using machine learning and contemporary deep learning architectures in analyzing big data in economics Mullainathan & Spiess (2017). For emerging market economies, the study of inflation forecasting using advanced ML models is in a nascent stage. Applications of neural networks are evident in developing countries like India, Brazil Araujo & Gaglianone (2023), Russia Pavlov (2020), Turkey, South Africa Botha et al. (2023), and China Özgür & Akkoç (2021). However, none of the above studies, to the best of our knowledge, have considered exogenous factors like macroeconomic uncertainties for inflation forecasting, especially for emerging markets.
@@
Following events such as the global financial crisis, it has been observed that many countries worldwide have encountered heightened uncertainty. This has had an impact on the decision-making process of individuals and consequently on the overall macroeconomy Bloom (2009, 2014); Baker et al. (2016). Although previous studies have primarily examined the connection between uncertainty and the overall economy, the evaluation of the changing relationship between measures of uncertainty and inflation is still in its early stages Jones & Olson (2013); Leduc & Liu (2016). Gaining insight into the intricacies of inflation dynamics and uncertainty metrics is critical for policymakers. Multiple recent studies indicate that both economic policy uncertainty and geopolitical risk significantly influence the movement of inflation within a given sample Colombo (2013); Jones & Olson (2013); Balcilar et al. (2017); Benchimol & El-Shagi (2020); Caldara & Iacoviello (2022); Adeosun et al. (2023); Anderl & Caporale (2023). Consequently, inflation forecasting models that do not incorporate these uncertainties are deemed to be incorrectly specified. The study conducted by Balcilar et al. (2017) examines the effectiveness of different uncertainty measures in predicting inflation. In particular, they use US inflation data and the economic policy uncertainty index (EPU) to compare how well the vector autoregressive fractionally integrated moving average (VARFIMA) model performs at making predictions compared to other models. The study conducted by Salisu et al. (2023) examined the predictive impact of EPU on stock market volatility in the United States and the United Kingdom. The findings suggest that EPU has a significant positive effect on volatility. However, this positive relationship weakens when the signal quality is low. Liu & Zhang (2015) asserted that EPU has a crucial impact on enhancing the predictive accuracy of models forecasting stock market volatility. The study conducted by Jones & Sackley (2016) investigated the impact of policy uncertainty in the United States and Europe on the short-term pricing model for gold. The study conducted by Junttila & Vataja (2018) investigated the impact of EPU on predicting future real economic activities in the UK and European economies. For an extensive analysis and thorough examination of different methods for quantifying uncertainty, as well as the impact of economic policy uncertainty on various economic activities such as stock market returns, corporate capital investments and spending, corporate finance, and risk management Al-Thaqeb & Algharabali (2019). On the other hand, the effect of geopolitical risk (GPRC) has been studied extensively by Caldara & Iacoviello (2022). A recent study by Pringpong et al. (2023) aims to link geopolitical risk with the reduction of firm values as well. The studies covering the impact of various uncertainty measures on different economic activities are still in the nascent phase, and this study, in turn, presents a way to accommodate the roles of EPU and GPRC in forecasting inflation numbers.
@@
Wavelet analysis has shown remarkable promise in analyzing financial time series due to its ability to distillate crucial information from the data. Their structural flexibility allows them to handle irregular time series data more effectively for long-term time series forecasting Percival & Guttorp (1994). Broadly, wavelets decompose the original time series into two sub-components: one concerning the low-frequency or deterministic component and the other concerning the high-frequency or stochastic component of the series. This process aims to reduce the effect of ‘noise’ in the forecast and is found to be quite useful in dealing with a variety of non-stationary signals, a common feature of financial and macroeconomic time series data, as they are constructed over finite time intervals and can represent the ‘local’ elements in both time and scale Percival & Walden (2000). Several applications of wavelet analysis can be found in forecasting exchange rates, GDP growth, crude oil prices, stock market prices, Taiwan stock index, S&P dividend yield, trade prices, expenditure and income, price movements, money growth, and inflation Yousefi et al. (2005); Shrestha & Tan (2005), volatility in foreign exchange markets, car sales from the last decades Crowley (2007). A wide variety of hybrid models combining wavelets and neural networks have been studied for forecasting natural gas prices Jin & Kim (2015), future trading Zhang et al. (2001), short-term electricity load Zhang & Dong (2001); Benaouda et al. (2006), energy price forecasting Tan et al. (2010), oil prices Yousefi et al. (2005), stock index Dunis & Williams (2002). Several studies have employed wavelet-based forecasting methods in various other domains of economics, business, and finance. For instance, Fernandez (2007) focused on the U.S. metal and materials manufacturing industry, employing wavelet decomposition combined with support vector machines to enhance forecasting accuracy. While the study effectively captures complex dynamics in manufacturing, it does not address long-term forecasting horizons or the inclusion of exogenous variables, which are crucial for economic forecasting. Rua (2017) used wavelet-based multiscale principal component analysis to improve factor model estimation and forecasting of macroeconomic variables like GDP growth and inflation in the U.S. Another paper by Souropanis & Vivian (2023) enhanced realized volatility (RV) forecasts for the S&P 500 using wavelet decomposition coupled with autoregressive model. However, these approaches do not address the non-linearity issue, as evident in CPI inflation data. In recent work, He et al. (2021) used a deep learning framework enhanced by wavelet transform to predict multivariate time series. Although their model captures the long-term and short-term characteristics of time series in the temporal domain through convolutional neural networks (CNN), it cannot be generalized well for limited inflation time series data problems. Vogl et al. (2022) utilized the efficacy of combining wavelet functions with neural networks for financial market predictions. More recently, Boubaker et al. (2023) proposed a hybrid model based on ARFIMA and wavelet neural network for forecasting the DJIA Index. However, the exclusion of exogenous variables and focus on high-frequency data limits its applicability to macroeconomic forecasting of CPI inflation. Apart from forecasting applications, wavelet analysis has been used in various areas of finance: Kim & In (2005) analyzed the association between stock price and inflation rate using wavelet analysis; Rua & Nunes (2012) explored the aspect of assessment of the risk in the emerging markets; He et al. (2012) studied the value at risk in the metal markets. Alexandridis & Hasan (2020) examined the impact of the global financial crisis on the multi-horizon nature of systematic risk and market risk using a wavelet multi-scale approach. Becerra et al. (2005) used wavelet neural networks for the prediction of corporate financial distress. However, none of the above studies have utilized the inflation series and its causal counterparts for long-term forecasting.
@@
Against this background, the primary aim of this study is to create a filtered ensemble wavelet neural network (FEWNet) model for CPI inflation forecasting. The key objective here is to generate accurate long-term forecasts for CPI inflation in BRIC countries while taking into account EPU and GPRC. FEWNet’s ability to handle non-linear, non-stationary, and complex time series further underscores its robustness. Its relatively stronger performance against baselines ranging from classical time series to modern deep learning architectures establishes its superiority in inflation forecasting contexts. Additionally, macroeconomic policy variables, characterized by low-frequency data, often present challenges where methods effective for high-frequency data (particularly various neural network architectures) perform sub-optimally. In analyzing macroeconomic policy variables like CPI inflation, the wavelet decomposition can extract the low-frequency components of the series without sacrificing the fundamental characteristics of the original time series. Our model also integrates a layer for feature engineering aspects by extracting the trend components using the Hodrick-Prescott (HP) filter and cyclical elements using the Christiano-Fitzgerald (CF) ideal band-pass filter in the forecasting exercise. Initially, the observed time series (CPI inflation) and exogenous factors (EPU and GPRC) undergo decomposition into their trend and cyclical components using HP and CF filters in the first phase. Subsequently, in the second phase, all the trend and cyclical components are modeled as auxiliary inputs within the FEWNet framework, along with the wavelet-transformed details and smooth coefficients of the CPI inflation. This comprehensive approach aims to generate forecasts for 12 months and 24 months ahead. Conditioning inflation forecasts on policy uncertainty and geopolitical risk would capture the inflation or deflation effects of uncertainty. We argue using wavelet coherence analysis that there exists a relationship between EPU and GPRC with that of CPI inflation for BRIC economies. FEWNet generates out-of-sample forecasts in a recursive manner, and using several statistical metrics and tests, we compare its performance with alternative forecasting methods. Theoretical results on empirical risk minimization are also presented for the FEWNet. Our study evaluates a wide array of neural network architectures, along with several statistical benchmarks, and demonstrates that FEWNet delivers superior forecasting performance for both semi-long-term (12 months ahead) and long-term (24 months ahead) forecast horizons for CPI inflation series in the BRIC countries. Apart from improvement in forecast accuracy measures of CPI inflation for the BRIC countries, another critical area that interests the policymakers is the knowledge of how various forms of structural uncertainties can influence the movement or complex dynamic behavior of different macroeconomic variables. The performance of FEWNet is superior to its counterparts in all three economies, except for China. This demonstrates that FEWNet is highly effective in handling complicated, non-stationary time series in forecasting tasks.
@@
The paper is organized in the following way. Section 2 explains the macroeconomic series considered in this study, various data pre-processing strategies, and the causal association between the uncertainty indices and CPI inflation. Section 3 focuses on the proposed methodology with a detailed overview of its architectural design, and Section 4 theoretically establishes the robustness of the method from an empirical risk minimization perspective. Section 5 discusses the experimental evaluation and the process of deriving a conformal prediction interval for the FEWNet framework. We discuss the policy implications this paper leads to in Section 6. Finally, Section 7 ultimately wraps up the study by presenting a thorough analysis of the main discoveries and outlining potential future directions.